OpenStudy (anonymous):

Simplify the expressions. Write your answer in lowest terms and leave it in factored form.

OpenStudy (anonymous):

\[\frac{ 2y }{y(3y-1) }+\frac{ 3 }{3y-1 }=\]

OpenStudy (anonymous):

@Jhannybean can you help me please

OpenStudy (anonymous):

this is what I am getting\[\frac{ 4 }{3y-1 }\]